Future Limited has the option to invest in Project X and Project Y but finance is only available to invest in one of them. You are given the following projected data: Project X Project Y Investment R900 000 R600 000 Depreciation method Straight-line Straight-line Cost of capital 14% 14% Economic life span 5 years 5 years Residual value at end of term R84 000 Nil Net profit: Year 1 56 000 70 000 Year 2 80 000 70 000 Year 3 102 000 70 000 Year 4 130 000 70 000 Year 5 152 000 70 000

Required:

3.1 Calculate the accounting rate of return for Project X. (Use average investment) (Answer must be expressed to two decimal places)

3.2 Calculate the payback period for both the projects. (Answer must be expressed in years, months and days)

3.3 Calculate the net present value for both the projects. (Round off amounts to the nearest Rand)

3.4 Which project would you choose and why?
Answer #1
Project
Year
Cashflow of Project X Cashflow of Project Y PV @ 14% Cost of Capital Cashflow of Project X Cashflow of Project Y
Year 0    (900,000.00) (600,000.00) 1.000 (900,000.00) (600,000.00)
Year 1      319,200.00    190,000.00 0.877    280,000.00    166,666.67
Year 2      243,200.00    190,000.00 0.769    187,134.50    146,198.83
Year 3      265,200.00    190,000.00 0.675    179,002.45    128,244.59
Year 4      293,200.00    190,000.00 0.592    173,597.94    112,495.25
Year 5      315,200.00    190,000.00 0.519    163,705.00      98,680.05
Year 5        84,000.00                   -   0.519      43,626.97                   -  
Ans 3.3) Net Present Value         127,067           52,285
Ans 3.1) ARR = Avg Annaul Profit / Avg Investment 25% 23%
Ans 3.2) Payback Period 3 Years 2 Months 29 Days 3 Years 1 Month 27 Days
Cash Inflow = Profit + Depreciation
Profit:
X      156,000.00      80,000.00 102,000.00    130,000.00    152,000.00 620,000.00
Y        70,000.00      70,000.00     70,000.00      70,000.00      70,000.00 350,000.00
Avg Annual Profit = Profit of all years / No of years
Avg Annual Profit Project X 124,000.00
Project Y     70,000.00
Average Investment = (Initial Investment + Salvage Value)/2
Average Investment Project X 492,000.00
Project Y 300,000.00
ARR Project X 25%
Project Y 23%
Formula for PV = 1/(1+Rate)^n For First Year = 1/(1+0.14)^1 = 0.877

Ans 3.4) The project with the best Net Present value should be chosen. In this case, it is Project X with higher NPV and the differential NPV is R 74781 which is 1.43 times of the NPV pf Project Y.